Minutes of the Tracker Meeting

October 28, 2004


Action Items:

  1. Nanda: find out what percentage of flight closeouts have been machined already by Plyform.
  2. Sandro: approve Jeff’s pro-forma for the prepreg shipment.
  3. Steve: look for an off-site vendor for drop testing.
  4. Steve: prepare the inner container proof loading plan for off-site execution.
  5. Jeff: procure 4-40 screws of the correct steel


0) Ben said that GSFC is busy processing tiles.  He did not know about the test status in Italy, so Robert reported that last week the tray with Goddard tiles completed testing up to 70C with no problems.


1) G&A will have the first two trays assembled tomorrow, and they will be acceptance tested on Tuesday.  Pisa is reworking the face-sheet/bias-circuit trimming on some trays, which were trimmed with too large a radius is some inside corners, causing interference with MCMs.


2) 21 completed tray panels have been delivered by Plyform, and they are either in Rome, at G&A, or in Pisa.  Robert asked what fraction of the flight closeouts have been machined.  Nanda will find out.


3) The Tower-A sidewalls are in Pisa.  The latest prepreg shipment is still in customs in Rome, as far as anybody knows.  INFN Pisa needs to review the pro-forma of the next (and last) shipment before Jeff releases it for shipping.

- Update on Plyform tray-panel production; closeout machining & assembly; Nanda will find out


4) Tower-B MCMs need to go through a cleaning operation on the connectors to remove conformal coating.  This will start tomorrow.  Also, some touch-up of conformal coating on the board may be needed on some of them.  Then the will get a repeat functional test.  The plan is for Persis to hand carry them on Nov 6.


5) Flex-circuit cables for Tower-A.  Ester is at Parlex doing source inspections and finding lots more quality issues.  Parlex is ready to ship the following cables (with the via problems corrected): 2-C2, 1-C5, 1-C6.  It is not looking very good to get new conforming cables in time for Tower A.  There is some discussion of completing the tower with some Tower-0 cables, doing environmental tests, and then replacing the non-conforming cables.  SLAC is increasingly frustrated with Parlex.  A second vendor has been found that claims to be able to make these cables, so a PR was put in for them to make some first articles.  SLAC is still pursuing a spliced prototype as well.


6) Other hardware deliveries for Tower-A:

a) Grid I/F hardware: a PO is established at Nyloc.  Arthur will go down Wednesday for source inspection.  SLAC should also ship more cones and hex bushings, but Jeff wants to get a full count from AMT first.

b) Peter said that one more week is needed to get socket wrenches and extraction tools.

c) Cables clips for the cable holding fixture are in hand and inspected.  They need a fit check with the new retainers.  ESD caps are ordered and expected in 1 more week; Jeff will check the status.  Screws are in hand but need a bakeout (Steve).

d) Corner brackets: 2 tower sets are complete.  They were 100% inspected by the vendor, but Jeff wants to validate the CMM measurements at SLAC. Clips are in, with the edges radiused.  Jeff ordered Kapton tape (1-inch and 2-inch widths), which is off the shelf.

d) The outer shipping container is at SLAC.  It still needs a drop test for qualification.  Steve will take care of this, but the SLAC stand-down is a problem.  Steve will look off-site for a vendor to do this.  It will ship to INFN with a dummy load and a monitor. 

e) Inner container: Diamond Tool is to complete it, since the SLAC shops are not working (it needs alodyne and inserts installed).  It needs assembly and proof load at SLAC.  Steve needs to formalize the procedure so that the proof load can be done out of house.  We still need the hoist rings.  Note that the Tower-0 inner container in Italy is usable for Tower-A.

f) Monitors: Pisa has 9 or 10 monitors and SLAC has 1.

g) Tower base plate assembly: 10 units have been received and the CMM done.  The grounding screw modification will have to be done out of house.

h) Arthur said that the cable bending tools can be done next week for both types, 1 of each.  Then they will do another 4 of each.

i) Cable holding plate: 20 units are on hand.  It will be proof loaded and travel along with the inner container.

j) Cap screws will be in this week or next week for the corner brackets. 

k) M4 8mm screws were delivered but need inspection.

l) Nicola reported that the 4-40 screws in Italy for the thermal straps are labeled 316 steel but should be A286.  Jeff will look into correcting this.

m) 100-degree screws should arrive today at SLAC from Italy.  Jeff is already working with a vendor that can give a 1-week lead on new screws. 

n) Heat strap tie downs are done but need inspection.

o) Solithane is due in today.  It is not hazmat, so it will ship to Pisa in the normal way.

p) ESD wrap is on order, 1 week out.

q) An entire drum of desiccant was ordered, 1 week out.

r) Jeff will order Poron foam today in several varieties.  The idea is to put the transfer adhesive on one side and stick it onto the sidewalls, to push the cables up against the closeouts.  It has 1 week lead time.  A test of it needs to be done in a vacuum chamber.

s) Nusil adhesives should be in Pisa; Macros says that they are in customs.

t) Torque wrenches: Pisa will update the list with the current status, quantities, and re-cal date and what SLAC needs to do.

u) Ball nests are being assembled at SLAC. 


7) Tower Assembly MRR: a tentative date is Nov 10.  This will be finalized next Thursday.